# SegAssess: Panoramic quality mapping for robust and transferable unsupervised segmentation assessment

### v1

@@ VI Limitations and future work
While SegAssess demonstrates impressive generalization across numerous datasets and achieves reliable zero-shot transferability to unseen segmentation masks from both model and human sources, certain limitations warrant further investigation. Although significantly improved compared to prior methods, SegAssess can still exhibit confusion between closely related PQM categories (i.e., intra-wrongness confusion between FP and FN, and intra-correctness confusion between TP and TN) in particularly challenging scenarios (Fig. 15). Furthermore, our ablation study on backbone sizes (see Section V-A1) indicated diminishing returns with larger models (SAM-L, SAM-H), suggesting that the current scale of available SQA training data may limit the effective utilization of these larger architectures.Finally, it is important to clarify the scope of the demonstrated zero-shot capability: SegAssess excels at assessing unseen masks within the learned domain (e.g., buildings, roads, water in high-resolution optical imagery), but its ability to generalize in a zero-shot manner to entirely new image domains (e.g., different sensor types) or fundamentally different object types without retraining has not been established and likely remains limited.

### v2

@@ VI Limitations and future work
While SegAssess demonstrates impressive generalization across numerous datasets and achieves reliable zero-shot transferability to unseen segmentation masks from both model and human sources, certain limitations warrant further investigation. Although significantly improved compared to prior methods, SegAssess can still exhibit confusion between closely related PQM categories (i.e., intra-wrongness confusion between FP and FN, and intra-correctness confusion between TP and TN) in particularly challenging scenarios (Fig. 15). Furthermore, our ablation study on backbone sizes (see Section V-A1) indicated diminishing returns with larger models (SAM-L, SAM-H), suggesting that the current scale of available SQA training data may limit the effective utilization of these larger architectures.Finally, Section V-D clarify the scope of the demonstrated zero-shot capability: SegAssess excels at assessing unseen masks within the learned image and object domain (e.g., buildings, roads, water in high-resolution optical imagery), but its ability to generalize in a zero-shot manner to entirely new image domains or fundamentally different object types remains limited.